\begin{thebibliography}{99}

\bibitem[DeLu]{DeLu} P. Deligne, G. Lusztig,
``Representations of reductive groups over a finite field",
Annals Math. {\bf 103} (1976) 103--161

\bibitem[DiMi]{DiMi} F. Digne, J. Michel,
\emph{Representations of Finite Groups of Lie Type 2nd ed.},
London Mathematical Society Student Texts {\bf 95}, 
Cambridge University Press, 2020

\bibitem[Fuj]{Fuj} A. Fujii,
``Parametrization of supercuspidal representations of depth zero for simple adjoint groups",
arXiv:2504.17225, 2025

\bibitem[GeMa]{GeMa} M. Geck, G. Malle,
\emph{The character theory of finite groups of Lie type. A guided tour},
Cambridge studies in advanced mathematics {\bf 187}, Cambridge University Press, 2020

\bibitem[Lus1]{Lus1} G. Lusztig,
\emph{Characters of reductive groups over a finite field},
Annals of Mathematics Studes {\bf 107},
Princeton University Press, Princeton NJ, 1984

\bibitem[Lus2]{Lus2} G. Lusztig,
``On the representations of reductive groups with disconnected centre",
Ast\'erisque {\bf 168} (1988), 157--166  

\bibitem[LuYu1]{LuYu1} G. Lusztig, Z. Yun,
``Endoscopy for Hecke categories, character sheaves and representations",
Forum Math. Pi {\bf 8} (2020), doi.org/10.1017/fmp.2020.9

\bibitem[LuYu2]{LuYu} G. Lusztig, Z. Yun
``Corrigendum to ‘Endoscopy for Hecke categories, character sheaves and representations´ "
Forum Math. Pi {\bf 9} (2021), doi.org/10.1017/fmp.2021.14 

\bibitem[MoPr]{MoPr} A. Moy, G. Prasad,
``Jacquet functors and unrefined minimal K-types",
Comment. Math. Helvetici {\bf 71} (1996), 98-121

\bibitem[Spr]{Spr} T.A. Springer,
\emph{Linear algebraic groups 2nd ed.},
Progress in Mathematics {\bf 9}, Birkh\"auser, Boston MA, 1998

\bibitem[Ste]{Ste} R. Steinberg,	
``Endomorphisms of linear algebraic groups",
Mem. Amer. Math. Soc. {\bf 80}, American Mathematical Society, Providence RI, 1968

\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{latexsym}
\usepackage{graphicx}
\usepackage{hyperref}
\usepackage{enumerate}
\usepackage[all]{xy}
\usepackage{chngcntr}

\newtheorem{thm}{Theorem}[section]

\newtheorem{lem}[thm]{Lemma}

\newtheorem{thmintro}{Theorem}

\theoremstyle{definition}

\newtheorem{cond}[thm]{Condition}

\newcommand{\Q}{\mathbb Q}

\newcommand{\F}{\mathbb F}
\newcommand{\mf}{\mathfrak}
\newcommand{\mc}{\mathcal}
\newcommand{\mb}{\mathbf}

\def\Irr{{\rm Irr}}
\newcommand{\mr}{\mathrm}

\newcommand{\Fr}{\mathrm{Fr}}

\newcommand{\Rep}{\mathrm{Rep}}
\newcommand{\Res}{\mathrm{Res}}

\newcommand{\der}{\mathrm{der}}

\begin{document}

\title[Representations of disconnected reductive groups over finite fields]
{Endoscopy for representations of disconnected reductive groups over finite fields}

\author{Maarten Solleveld}
\address{IMAPP, Radboud Universiteit Nijmegen, Heyendaalseweg 135,
6525AJ Nij\-megen, the Netherlands}
\email{m.solleveld@science.ru.nl}
\subjclass[2010]{20C33, 20G40}

\maketitle

\begin{abstract}
Let $G$ be the group of rational points of a connected reductive group over a finite field.
Based on work of Lusztig and Yun, we make the Jordan decomposition for irreducible 
$G$-representations canonical. It comes in the form of an equivalence between the category of
$G$-representations with a fixed semisimple parameter $s$ and the category of unipotent 
representations of an endoscopic group of $G$, enriched with an equivariant structure
with respect to the component group of the centralizer of $s$.

Next we generalize these results, replacing the connected reductive group by a 
smooth group scheme with reductive neutral component. Again we esta\-blish canonical
equivalences for both the rational and the geometric series of $G$-representations, 
in terms of unipotent representations of endoscopic groups.
\end{abstract}

\section{Introduction}

Let $\mc G$ be a connected reductive group over a finite field $\F_q$. We are interested 
in the representations of $G = \mc G (\F_q)$, always on vector spaces over $\overline{\Q}_\ell$
where $q$ and $\ell$ are coprime.

Deligne and Lusztig \cite{DeLu} showed that the irreducible $G$-representations can be 
partitioned in series, indexed by semisimple conjugacy classes in the $\F_q$-group $\mc G^\vee$
dual to $\mc G$. More concretely, starting from a maximal $\F_q$-torus $\mc T$ in $\mc G$ and
a character $\theta$ of $T = \mc T(\F_q)$ they constructed a virtual $G$-representation $R_T^G (\theta)$. 
The data $(\mc T,\theta)$ determine a semisimple element $s \in G^\vee = \mc G^\vee (\F_q)$,
up to $G^\vee$-conjugacy. Moreover, it was shown in \cite{DeLu} that $R_T^G (\theta)$ and 
$R_{T'}^G (\theta')$ have no common irreducible constituents if their semisimple parameters 
not geometrically conjugate. That enables the following definition: the geometric series with 
parameter $s$ consists of the irreducible $G$-representations that appear in 
$R_{T}^G (\theta)$ for some $(\mc T, \theta)$ whose associated element in $sG^\vee$ 
is geometrically conjugate to $s$. 

Later Lusztig \cite{Lus1} improved on that, by showing that two virtual representations 
$R_T^G (\theta)$ and $R_{T'}^G (\theta')$ are disjoint whenever their parameters in $G^\vee$ 
are not $G^\vee$-conjugate. This lead to the rational (or Lusztig) series $\Irr_s (G)$, where $s$
represents a semisimple (rational) conjugacy class in $G^\vee (\F_q)$. The element $s$ is called
the semisimple parameter of a representation in $\Irr_s (G)$, uniquely determined up to
$G^\vee$-conjugacy. The representations in $\Irr_1 (G)$ are especially interesting, these are
known as unipotent representations. 

In \cite{Lus1,Lus2}, Lusztig proved that there is always a bijection
\begin{equation}\label{eq:A.1} 
\Irr_s (G) \longleftrightarrow \Irr_1 (Z_{G^\vee}(s)) .
\end{equation}
This works best when $Z_{\mc G^\vee}(s)$ is connected, then the bijection is canonical. Thus 
every irreducible $G$-representation is determined by its semisimple parameter and a unipotent
representation of another reductive $\F_q$-group. This is known as the Jordan decomposition of
irreducible $G$-representations, we refer to \cite{DiMi,GeMa} for much more background.\\

In \cite{LuYu1,LuYu}, Lusztig and Yun set out to make \eqref{eq:A.1} canonical, in geometric
terms. This required several modifications. Firstly, they do not work with rational but with
geometric series. A maximal $\F_q$-torus $\mc T$ in $\mc G$ is fixed, say with Weyl
group $W$. Characters of $T$ are replaced by rank one character sheaves on $\mc T$. These are
not required to be stable under the action of a Frobenius element $\Fr$, only the 
$W$-orbit of $\mc L$ must be $\Fr$-stable. Such data $(W\mc L,\mc T)$ correspond to precisely
one $\Fr$-stable geometric conjugacy class in $G^\vee$.

Secondly, the main results of Lusztig--Yun involve categories, not only irreducible 
representations. Instead of $\Irr_x (G)$ (given some data $x$ for which this is defined) we have 
$\Rep_x (G)$, the category generated by $\Irr_x (G)$.
In particular $\Rep_{W \mc L}(G)$ denotes the geometric series associated to $W\mc L$.

Thirdly, $Z_{\mc G^\vee}(s)$ is replaced by the endoscopic group $\mc H$ dual to 
$Z_{\mc G^\vee}(s)^\circ$ plus some group $\Omega_{\mc L}$ isomorphic to $\pi_0 (Z_{\mc G}(s))$.
In this way the use of $\mc G^\vee$ can be avoided in \cite{LuYu1,LuYu}.

With impressive geometric machinery Lusztig and Yun obtain canonical equivalences between
subcategories of $\Rep_{W\mc L}(G)$ (determined by cells) and subcategories of $\Rep_1 (H)$
enriched with $\Omega_{\mc L}$-equivariant structures. This can be regarded as an endoscopic
categorical version of Deligne--Lusztig theory.\\

The aims of this paper are threefold. Firstly, we want to explain the final results of 
\cite{LuYu} in simpler representation theoretic terms. Secondly, we want to find a version
of these results for rational series, so a canonical Jordan decomposition of irreducible
$G$-representations, like \eqref{eq:A.1}. This may have been known to Lusztig and Yun, but it
is not stated in \cite{LuYu1,LuYu}. And thirdly, we want to generalize these canonical 
equivalences of categories to disconnected reductive $\F_q$-groups. In a sense, this is the
natural generality for such results. Namely, \eqref{eq:A.1} already shows that to analyse
irreducible representations of connected reductive $\F_q$-groups, one has to involve unipotent 
representations of disconnected reductive $\F_q$-groups. If instead we start with a disconnected 
reductive $\F_q$-group, then we still need to study unipotent representations of other 
disconnected reductive $\F_q$-groups, but we stay within the same class of groups.\\

Now we discuss our main results in more detail. Let $\mc G$ be a smooth $\F_q$-group scheme
whose neutral component $\mc G^\circ$ is reductive. Only the connected components of $\mc G$
which are stable under a Frobenius element $\Fr$ contribute to $G = \mc G (\F_q)$, so we may 
assume that all connected components of $\mc G$ are $\Fr$-stable. 
We allow $\pi_0 ( \mc G)$ to be infinite.

Let $\mc L$ be a rank one character sheaf on a maximal $\F_q$-torus $\mc T$ in $\mc G^\circ$,
such that $W(\mc G^\circ,\mc T) \mc L$ is $\Fr$-stable. We define the geometric series 
$\Rep_{\mc W (\mc G,\mc T) \mc L}(G)$ as consisting of
the $G$-representations $\pi$ such that $\Res^G_{G^\circ} \pi$ lies in the subcategory generated
by $\bigcup_{w \in W(\mc G,\mc T)} \Rep_{W(\mc G^\circ,\mc T) w \mc L} (G^\circ)$.

Let $\mc H$ be the (connected) endoscopic group of $\mc G^\circ$ determined by $(\mc L,\mc T)$. 
To $(\mc L,\mc T)$ we also associate a group $\Omega_{\mc L}$, which plays the role of the component 
group of some disconnected version of $\mc H$. 
We need a certain set $\mc B_{\mc L}^\circ$ with an action of $\Omega_{\mc L}^\circ$ (the version
of $\Omega_{\mc L}$ for $\mc G^\circ$) such that $\mf B_{\mc L}^\circ / \Omega_{\mc L}^\circ$
parametrizes the rational conjugacy classes in one geometric conjugacy class in $G^{0 \vee}$.
Every $\beta \in \mf B_{\mc L}^\circ$ gives rise to a $\F_q$-form of $\mc H$, say $\mc H^\beta$.

\begin{thmintro}\label{thm:A}
There exists a canonical equivalence of categories
\[
\Rep_{W(\mc G,\mc T) \mc L} (G) \cong \Big( \bigoplus\nolimits_{\beta \in \mf B_{\mc L}^\circ}
\Rep_1 \big( H^\beta \big) \Big)^{\Omega_{\mc L}} .
\]
\end{thmintro}

The superscript $\Omega_{\mc L}$ means that we enrich the objects in the category between 
brackets with $\Omega_{\mc L}$-equivariant structures. Thus Theorem \ref{thm:A} describes 
$G$-represen\-ta\-tions with semisimple geometric parameter $W(\mc G,\mc T) \mc L$ in terms of 
unipotent representations of forms of an endoscopic group $\mc H$ plus a kind of action of
the group $\Omega_{\mc L}$. When $\mc G$ is connected, Theorem \ref{thm:A} is a quick consequence 
of \cite[Corollary 12.7]{LuYu}.

Suppose now that $\mc L$ is $\Fr$-stable. Then it determines a semisimple rational conjugacy
class in $G^{\circ \vee}$, say represented by $s$. We define the rational series
$\Rep_s (G)$ as the category of $G$-representations $\pi$ such that $\Res^G_{G^\circ} \pi$ 
lies in the subcategory generated by $\bigcup_{g \in G} \mr{Ad}(g)^* \Rep_s (G^\circ)$.

\begin{thmintro}\label{thm:B}
There exists a canonical equivalence of categories
\[
\Rep_s (G) \cong \Rep_1 (H)^{\Omega_{\mc L}^\Fr} .
\]
\end{thmintro}

This can be regarded as a Jordan decomposition for representations of the possibly disconnected
reductive group $G = \mc G (\F_q)$. It existed already when $\mc G$ is connected \cite{Lus2},
but even in that case Theorem \ref{thm:B} adds canonicity.\\

We expect that our results will have applications to depth zero representations of reductive
$p$-adic groups. The main reason is that by \cite{MoPr} every supercuspidal depth zero
representation of such a group, say $M$, is obtained from an irreducible representation of
a group of the form $\mc G (\F_q)$. In that setting $\mc G (\F_q)$ arises as a quotient of 
the $M$-stabilizer of a vertex in the reduced Bruhat--Tits building of $M$.

In the recent preprint \cite{Fuj}, depth zero supercuspidal representations of simple $p$-adic
groups were analysed, with several techniques among which some version of a Jordan decomposition
for $\Irr (\mc G (\F_q))$. With Theorem \ref{thm:B} it should be possible to make the results in 
\cite[\S 4]{Fuj} applicable in larger generality.

\renewcommand{\theequation}{\arabic{section}.\arabic{equation}}
\numberwithin{equation}{section}

\section{Connected groups} 

We start by recalling a part of the setup and the relevant results from \cite{LuYu1,LuYu}.
Let $\mc G$ be a connected reductive group defined over the algebraic closure of a finite
field $\F_q$. Let $\ell$ be a prime number different from $p = \mr{char}(\F_q)$. All our
representations will be on vector spaces over $\overline{\Q_\ell}$.

Let $\epsilon$ be the action of a geometric Frobenius element $\Fr \in 
\mr{Gal}(\overline{\F_q} / \F_q)$ for some rational structure of $\mc G$. We denote the 
corresponding $\F_q$-group by $\mc G^\epsilon$, and we write
\[
G^\epsilon = \mc G^\epsilon (\F_q) = \mc G \big( \overline{\F_q} \big)^\epsilon .
\]
We fix an $\epsilon$-stable maximal torus $\mc T$ in $\mc G$, so $\mc T^\epsilon$ is a maximal
$\F_q$-torus in $\mc G^\epsilon$. Let $\mc B$ be an $\epsilon$-stable Borel subgroup of $\mc G$
containing $\mc T$ and let $\Phi^+$ be the corresponding set of positive roots in
$\Phi = \Phi (\mc G,\mc T)$. We write $W = W(\mc G,\mc T)$ and for each $w \in W$ we fix
a representative $\dot{w} \in N_{\mc G}(\mc T)$. 

Let $\mc L$ be a rank one character sheaf on $\mc T$, and assume that $W \mc L$ is 
$\epsilon$-stable. Whenever $w \epsilon \mc L = \mc L$, $\mc L$ can be regarded as a character 
sheaf on $\mc T^{w \epsilon}$, stable under the Frobenius action $\mr{Ad}(w) \circ \epsilon$ of
that $\F_q$-torus. Then $\mc L$ is equivalent to the data of a character
$T^{w \epsilon} \to \overline{\Q_\ell}^\times$.

The orbit $W \mc L$ corresponds to a Fr-stable semisimple conjugacy class in the 
reductive $\F_q$-group $\mc G^\vee$ dual to $\mc G$ \cite[\S 5]{DeLu}.
Deligne--Lusztig theory yields the category $\Rep_{W \mc L}(G^\epsilon)$ generated
by irreducible $G^\epsilon$-representations with semisimple parameter $W \mc L$. 
The set of irreducible representations in $\Rep_{W \mc L}(G^\epsilon)$ is 
called the geometric series with parameter $\mc L$ or $W \mc L$.

The stabilizer of $\mc L$ in $W$ can be written as 
\[
W_{\mc L} = W_{\mc L}^\circ \rtimes \Omega_{\mc L} .
\]
Here $W_{\mc L}^\circ$ is the Weyl group of the root system 
\[
\Phi_{\mc L} = \{ \alpha \in \Phi : (\alpha^\vee)^* \mc L \text{ is trivial} \}
\] 
and $\Omega_{\mc L}$ is the $W_{\mc L}$-stabilizer of $\Phi_{\mc L}^+ = \Phi^+ \cap \Phi_{\mc L}$.
Let $\mc H$ be the connected reductive $\overline{\F_q}$-group with maximal torus $\mc T$ and 
root system $\Phi_{\mc L}$. It is called the endoscopic group of $\mc G$ associated to $\mc L$. 
It can be endowed with a pinning relative to $\mc G$, as in \cite[\S 10.2]{LuYu}.
Consider the set
\[
\mf B_{\mc L} = W_{\mc L}^\circ \, \backslash \{ w \in W : w \epsilon \mc L = \mc L \} .
\]
The group $\Omega_{\mc L} \cong W_{\mc L}^\circ \, \backslash W_{\mc L}$ acts on $\mc B_{\mc L}$ by 
\begin{equation}\label{eq:E.26}
\mr{Ad}_\epsilon (\omega) w = \omega w \epsilon (\omega)^{-1}.
\end{equation}
Any other $\epsilon$-stable maximal torus in $\mc G$ has the form $g \mc T g^{-1}$ for some $g \in 
\mc G$. There are bijections
\[
\begin{array}{cccccc}
W_{\mc L} & \to & W_{g \mc L} \, : & w & \mapsto & g w g^{-1} , \\
\mf B_{\mc L} & \to & \mf B_{g \mc L} \, : & w & \mapsto & g w \epsilon (g)^{-1} ,
\end{array}
\]
which induce a bijection 
\[
\mf B_{\mc L} / \mr{Ad}_\epsilon (\Omega_{\mc L}) \to 
\mf B_{g \mc L} / \mr{Ad}_\epsilon (\Omega_{g \mc L}) .
\]
The data $(\mc L, \mc T, \beta)$ are considered equivalent to $(g \mc L, g \mc T g^{-1}, 
g \beta \epsilon (g)^{-1})$, and all the below constructions from these two sets of data related
by conjugation by $g$.

Each class $\beta \in \mf B_{\mc L}$ contains a unique element $w^\beta$ of minimal length,
which sends $\Phi_{\mc L}^+$ to $\Phi^+$. As explained in \cite[\S 12.1]{LuYu}, the relative
pinning and the lift $\dot{w}^\beta$ determine an automorphism $\sigma_{\beta \epsilon}$ of 
$(\mc H, \mc T, \Phi_{\mc L}^+)$, whose action on $\mc T$ and on the root subgroups in
$\mc H$ corresponds to the action of $\mr{Ad}(\dot{w}^\beta) \circ \epsilon$ on $\mc T$ 
and on the root subgroups in $\mc G$. This provides a Frobenius action on $\mc H$, and hence 
a $\F_q$-form $\mc H^{\sigma_{\beta \epsilon}}$ of $\mc H$. Up to a canonical isomorphism,
it does not depend on the choice of $\dot{w}^\beta$.

Next we want to define $\Omega_{\mc L}$-equivariant structures on $\bigoplus_{\beta \in 
\mf B_{\mc L}} \Rep_1 (H^{\sigma_{\beta \epsilon}})$. For every $\pi_\beta \in 
\Rep_1 (H^{\sigma_{\beta \epsilon}})$ and every $\omega \in \Omega_{\mc L}$ we have 
\[
\mr{Ad}(\dot \omega) \pi_\beta \in \Rep_1 \big( H^{\mr{Ad}(\dot \omega) \sigma_{\beta \epsilon} 
\mr{Ad}(\dot \omega)^{-1}} \big).
\]
For another representatives $\ddot{\omega}$, $\mr{Ad}(\ddot \omega) \pi_\beta$ differs from
$\mr{Ad}(\dot \omega) \pi_\beta$ by a canonical isomorphism, namely $\mr{Ad}(\ddot{\omega}
\dot{\omega}^{-1})$ with $\ddot{\omega} \dot{\omega}^{-1} \in \mc T$. Lang's theorem 
ensures the existence of an isomorphism
\[
H^{\mr{Ad}(\dot \omega) \sigma_{\beta \epsilon} \mr{Ad}(\dot \omega)^{-1}} \cong
H^{\sigma_{\mr{Ad}_\epsilon (\omega) \beta \epsilon}} .
\]
It is conjugation by an element of $\mc T$, which is unique up to $T^{\sigma_{\mr{Ad}_\epsilon 
(\omega) \beta \epsilon}}$. Via this isomorphism we regard $\mr{Ad}(\dot \omega) \pi_\beta$ 
as an element of $\Rep_1 (H^{\sigma_{\mr{Ad}_\epsilon (\omega) \beta \epsilon}})$.
Equivalently, this construction means that we select a representative $n(\omega,\beta) \in 
N_{\mc G}(\mc T)$ for $\omega$ (unique up $T^{\sigma_{\mr{Ad}_\epsilon (\omega) \beta \epsilon}}$
from the left and up to $T^{\sigma_{\beta \epsilon}}$ from the right), such that 
\[
\sigma_{\mr{Ad}_\epsilon (\omega) \beta \epsilon} = 
\mr{Ad}(n(\omega,\beta)) \sigma_{\beta \epsilon} \mr{Ad}(n(\omega,\beta))^{-1} .
\]
Then we define
\[
\mr{Ad}(\omega) \pi_\beta = \mr{Ad}(n(\omega,\beta)) \pi_\beta \in 
\Rep_1 (H^{\sigma_{\mr{Ad}_\epsilon (\omega) \beta \epsilon}}) .
\]
Here $\mr{Ad}(\omega) \pi$ is independent of the choices, up to canonical isomorphisms
(namely inner automorphisms, from $T^{\sigma_{\mr{Ad}_\epsilon (\omega) \beta \epsilon}}$).
That defines an action of $\Omega_{\mc L}$ on\\ $\bigoplus_{\beta \in 
\mf B_{\mc L}} \Rep_1 (H^{\sigma_{\beta \epsilon}})$ modulo inner automorphisms. 

Now an $\Omega_{\mc L}$-equivariant structure on $\pi = \bigoplus_{\beta \in \mf B_{\mc L}}
\pi_\beta$ consists of a morphism from $\pi$ to $\mr{Ad}(\omega)\pi$ for each
$\omega \in \Omega_{\mc L}$, multiplicative in $\omega$ up to canonical isomorphisms.
To make that explicit we need, for each representative $n(\omega,\beta)$, a morphism
\begin{equation}\label{eq:E.28}
\pi (n(\omega,\beta)) : \pi_\beta \to \mr{Ad}(n(\omega,\beta)) \pi_\beta 
\end{equation}
such that:
\begin{itemize}
\item $\pi (n(1,\beta)) = \pi_\beta (n(1,\beta))$ for all 
$n(1,\beta) \in T^{\sigma_{\beta \epsilon}}$,
\item for all eligible $n(\omega',\mr{Ad}_\epsilon (\omega) \beta)$:
\[
\pi \big( n (\omega',\mr{Ad}_\epsilon (\omega) \beta) n (\omega,\beta) \big) = 
\pi \big( n (\omega',\mr{Ad}_\epsilon (\omega) \beta) \big) \, \pi \big( n (\omega,\beta) \big) .
\]
\end{itemize}
The category $\bigoplus_{\beta \in \mf B_{\mc L}} \Rep_1 (H^{\sigma_{\beta \epsilon}})$
enriched with $\Omega_{\mc L}$-equivariant structures is denoted 
\[
\Big( \bigoplus\nolimits_{\beta \in \mf B_{\mc L}} 
\Rep_1 (H^{\sigma_{\beta \epsilon}}) \Big)^{\Omega_{\mc L}} .
\]
The following result is a version of \cite[Corollary 12.7]{LuYu} without cells.

\begin{thm}\label{thm:E.1}
Let $\mc G^\epsilon$ be a connected reductive $\F_q$-group and let $W \mc L$ be an 
$\epsilon$-stable orbit of character sheaves on $\mc T$. There is a canonical 
equivalence of categories
\[
\Rep_{W \mc L} (G^\epsilon) \cong \Big( \bigoplus\nolimits_{\beta \in \mf B_{\mc L}} 
\Rep_1 (H^{\sigma_{\beta \epsilon}}) \Big)^{\Omega_{\mc L}} .
\]
\end{thm}
\begin{proof}
Let $\mb c$ be a cell in $W_{\mc L}^\circ$ and let $[\mb c] \subset W \times W \mc L$ be
the associated two-sided cell, as in \cite[\S 11.4]{LuYu}. They are related by
\begin{equation}\label{eq:E.3}
[\mb c] \cap (W_{\mc L}^\circ \times \{\mc L\}) = \Omega_{\mc L} \mb c .
\end{equation}
By construction
\begin{equation}\label{eq:E.2}
\Rep_{W \mc L} (G^\epsilon) = \bigoplus\nolimits_{[\mb c]} \Rep_{W \mc L}^{[\mb c]} (G^\epsilon) . 
\end{equation}
Let $\Omega_{\mb c}$ be the stabilizer of $\mb c$ in $\Omega_{\mc L}$ and write
\[
\mf B_{\mb c} = \{ \beta \in \mf B_{\mc L} : w^\beta \epsilon \text{ stabilizes } \mb c \} .
\]
By \cite[Corollary 12.7]{LuYu} there are canonical equivalences of categories
\begin{equation}\label{eq:E.1}
\Rep_{W \mc L}^{[\mb c]} (G^\epsilon) \cong 
\bigoplus_{\beta \in \mf B_{\mb c} / \mr{Ad}_\epsilon (\Omega_{\mb c})} \Rep_1^{\mb c}
\big( H^{\sigma_{\beta \epsilon}} \big)^{\Omega_{\mb c,\beta}} \cong
\Big( \bigoplus_{\beta \in \mf B_{\mb c}} \Rep_1^{\mb c} \big( H^{\sigma_{\beta \epsilon}} \big)
\Big)^{\Omega_{\mc L}} .
\end{equation}
When $\beta \in \mf B_{\mc L} \setminus \mf B_{\mb c}$, $H^{\sigma_{\beta \epsilon}}$
has no nonzero representations associated to the cell $\mb c$ (because the cell of an
irreducible representation is unique). Hence the right hand side of \eqref{eq:E.1} can
be rewritten as
\[
\Big( \bigoplus\nolimits_{\beta \in \mf B_{\mc L}} \Rep_1^{\Omega_{\mc L} \mb c} 
\big( H^{\sigma_{\beta \epsilon}} \big) \Big)^{\Omega_{\mc L}} .
\]
Now we take the direct sum over all two-sided cells $[\mb c]$ and with \eqref{eq:E.2}
and \eqref{eq:E.3} we obtain
\[
\Rep_{W \mc L} (G^\epsilon) \cong \bigoplus_{[\mb c]} \Big( \bigoplus_{\beta \in 
\mf B_{\mc L}} \Rep_1^{\Omega_{\mc L} \mb c} \big( H^{\sigma_{\beta \epsilon}} \big) 
\Big)^{\Omega_{\mc L}} = \Big( \bigoplus_{\beta \in \mf B_{\mc L}} 
\Rep_1 \big( H^{\sigma_{\beta \epsilon}} \big) \Big)^{\Omega_{\mc L}} . \qedhere
\]
\end{proof}

We turn our attention to rational series for $G^\epsilon$. Let $\mc G^\vee$ be
the dual group of $\mc G$. It is endowed with the $\F_q$-structure from the Frobenius 
action dual to $\epsilon$, which we shall also denote $\epsilon$. 
Recall that every rational series is parametrized by the (rational) conjugacy 
class of a semisimple element $s$ of $G^{\vee \epsilon} = \mc G^\vee (\F_q)$. 
In terms of $(\mc L, \mc T^\epsilon)$, the condition $s \in G^{\vee \epsilon}$ means
that $\mc L$ must be $\epsilon$-stable. That can always be arranged:

\begin{lem}\label{lem:E.5}
Suppose that $W \mc L$ is $\epsilon$-stable. Then $(\mc L,\mc T)$ is $\mc G$-conjugate
to an $\epsilon$-stable pair $(\mc L_w, \mc T_w)$ and $\Rep_{W \mc L}(G^\epsilon) =
\Rep_{W(\mc G,\mc T_w) \mc L_w} (G^\epsilon)$.
\end{lem}
\begin{proof}
Select $w \in W$ such that $w \epsilon \mc L = \mc L$. By the surjectivity of Lang's map
on $\mc G$ \cite[Theorem 1.4.8]{GeMa}, there exists $g_w \in \mc G$ such that 
$g_w^{-1} \epsilon (g_w)$ represents $w \in N_{\mc G}(\mc T) / \mc T$. Then
\begin{align*}
& g_w \mc T g_w^{-1} = g_w w \mc T w^{-1} g_w^{-1} = \epsilon (g_w) \mc T \epsilon (g_W)^{-1} 
= \epsilon (g_w) \epsilon (\mc T) \epsilon (g_w)^{-1} = \epsilon (g_w \mc T g_w^{-1}) , \\
& g_w \mc L = g_w w \epsilon \mc L = \epsilon (g_w) \epsilon \mc L = \epsilon g_w \mc L , 
\end{align*}
so $(\mc L_w,\mc T_w) := (g_w \mc L, g_w \mc T g_w^{-1})$ is $\epsilon$-stable. The equality
of categories holds because $W \mc L$ and $W(\mc G,\mc T_w) \mc L_w$ correspond to the 
same geometric conjugacy in $G^{\vee \epsilon}$.
\end{proof}

With Lemma \ref{lem:E.5} in mind, we assume that $(\mc L,\mc T)$ is $\epsilon$-stable. 
Recall from \cite[\S 5.2]{DeLu} or \cite[Corollary 2.5.14]{GeMa} that there is a natural
bijection between 
\begin{itemize}
\item the set of $G^\epsilon$-conjugacy classes of $\epsilon$-stable pairs 
$(\mc L,\mc T)$,
\item the set of $G^{\vee \epsilon}$-conjugacy classes of $\epsilon$-stable
pairs in $G^{\vee \epsilon}$ formed by a maximal torus and an element thereof. 
\end{itemize}
We record that as
\begin{equation}\label{eq:E.22}
(\mc L,\mc T) \longleftrightarrow (s,\mc T^\vee).
\end{equation} 
The intersection of the geometric conjugacy class $\mr{Ad}(\mc G^\vee) s$ with $\mc T^\vee$ 
is $W s$ with $W \cong W(\mc G^\vee,\mc T^\vee)$. By \cite[\S 5]{DeLu} or \cite[Propositions 
2.4.28, 2.5.5 and p. 155]{GeMa}, this corresponds 
precisely to $(W \mc L,\mc T)$. Then $\Rep_{W \mc L}(G^\epsilon)$ is a direct sum of 
finitely many rational series, which we want to parametrize in terms of $\mc L$.

\begin{lem}\label{lem:E.6}
The quotient set $\mf B_{\mc L} / \mr{Ad}_\epsilon (\Omega_{\mc L})$ naturally parametrizes the 
$G^{\vee \epsilon}$-conju\-ga\-cy classes in $(\mr{Ad}(\mc G^\vee) s)^\epsilon$. 
The parametrization can be realized as follows. For $w \in W_{\mc L}$ representing 
$\beta \in \mf B_{\mc L}$, find $h_w \in \mc G^\vee$ such that $h_w^{-1} \epsilon (h_w)$ 
represents $w$ as element of $W \cong N_{\mc G^\vee}(\mc T^\vee) / \mc T^\vee$. Then the 
bijection sends $\beta = [w] \in \mf B_{\mc L} / \mr{Ad}_\epsilon (\Omega_{\mc L})$ to
$\mr{Ad}(G^{\vee \epsilon}) (h_w s h_w^{-1})$.
\end{lem} 
\begin{proof}
As $\Fr (s) = s$, the $\mc G^\vee$-conjugacy class of $s$ is isomorphic to 
$\mc G^\vee / Z_{\mc G^\vee}(s)$ as $\F_q$-variety. Galois cohomology provides an exact sequence
\begin{equation}\label{eq:E.11}
Z_{\mc G^\vee}(s)^\Fr \to (\mc G^\vee)^\Fr \to \big( \mc G^\vee / Z_{\mc G^\vee}(s) \big)^\Fr
\to H^1 (\Fr, Z_{\mc G^\vee}(s)) \to H^1 (\Fr ,\mc G^\vee) .
\end{equation}
Here $H^1 (\Fr ,\mc G^\vee)$ is a short notation for the cohomology of 
$\mr{Gal}(\overline{\F_q}/\F_q)$ with coefficients in $\mc G^\vee$. By the surjectivity of
Lang's map on connected $\F_q$-groups, $H^1 (\Fr ,\mc G^\vee)$ is trivial and 
\[
H^1 (\Fr, Z_{\mc G^\vee}(s)) \cong H^1 \big( \Fr, \pi_0 (Z_{\mc G^\vee}(s)) \big) . 
\]
Thus \eqref{eq:E.11} says that the set we are looking for is parametrized naturally by\\
$H^1 \big( \Fr, \pi_0 (Z_{\mc G^\vee}(s)) \big)$. This bijection sends 
\begin{equation}\label{eq:E.12}
g s g^{-1} \in (\mr{Ad}(\mc G^\vee) s)^\Fr \quad \text{to} \quad 
[\Fr \mapsto g^{-1} \Fr (g)] \in H^1 (\Fr, Z_{\mc G^\vee}(s)).
\end{equation}
By Steinberg's description of the centralizers of semisimple elements in reductive groups
\cite{Ste}, $\pi_0 (Z_{\mc G^\vee}(s)) \cong W_s^\circ \backslash W_s$, where $W_s^\circ
\subset W \cong W(\mc G^\vee,\mc T^\vee)$ is generated by the reflections $s_\alpha$ with
$\alpha^\vee (s) = 1$. In terms of $(\mc L,\mc T)$, this becomes 
\begin{equation}\label{eq:E.17}
\pi_0 (Z_{\mc G^\vee}(s)) \cong W_s^\circ \backslash W_s \cong 
W_{\mc L}^\circ \backslash W_{\mc L} = \mc B_{\mc L} = 
W_{\mc L}^\circ \backslash W_{\mc L}^\circ \rtimes \Omega_{\mc L} \cong \Omega_{\mc L}.
\end{equation}
It follows that $H^1 \big( \Fr, \pi_0 (Z_{\mc G^\vee}(s)) \big)$ is naturally isomorphic to 
\begin{equation}\label{eq:E.13}
H^1 (\Fr ,\Omega_{\mc L}) = (\Omega_{\mc L})_\Fr = 
\mf B_{\mc L} / \mr{Ad}_\epsilon (\Omega_{\mc L}) .
\end{equation}
The map from \eqref{eq:E.13} to conjugacy classes of $s$ is essentially the inverse of
\eqref{eq:E.12}.
\end{proof}

Lemmas \ref{lem:E.5} and \ref{lem:E.6} provide a bijection between 
\begin{itemize}
\item the set of semisimple conjugacy classes in $G^{\vee \epsilon}$,
\item the $W$-orbits of pairs $(\mc L,\beta)$, where $\mc L$ is a character sheaf on $\mc T$ 
with $\epsilon$-stable $W$-orbit and $\beta \in \mf B_{\mc L} / \mr{Ad}_\epsilon (\mc L)$,
\item triples $(\mc L', \mc T', \beta')$ where $\mc T'$ is an $\epsilon$-stable maximal torus
in $\mc G$, $(\mc L', \beta')$ is as in the previous bullet and any such triple of the form 
$(g \mc L', g \mc T' g^{-1}, g \beta' \epsilon (g)^{-1})$ with $g \in \mc G^\circ$ 
is considered as equivalent to $(\mc L', \mc T', \beta')$.
\end{itemize}
Let us make it more explicit, starting from $W (\mc L,\beta)$. As in Lemma \ref{lem:E.5},
pick $g \in \mc G$ such that $(g \mc L, g \mc T g^{-1})$ is $\epsilon$-stable, and
replace $\beta$ by $g \beta \epsilon (g)^{-1} \in \mf B_{g \mc L}$. Associate $s \in G^{\vee 
\epsilon}$ to $(g \mc L, g \mc T g^{-1})$ by \eqref{eq:E.22}. From $s$ and 
$g \beta \epsilon (g)^{-1}$, Lemma \ref{lem:E.6} produces $s_\beta = h_w s h_w^{-1}$,
where $w$ represents $g \beta \epsilon (g)^{-1}$. We record this as 
\begin{equation}\label{eq:E.23}
W (\mc L, \beta) \longleftrightarrow s_\beta .
\end{equation}
The explicit description of the correspondence 
\eqref{eq:E.22} in \cite[\S 2.5.12 and Corollary 2.4.14]{GeMa} is in terms of elements 
$h_w$ and $g_w$ as in the proof of Lemma \ref{lem:E.5}. Consequently 
\begin{equation}\label{eq:E.20}
(s_\beta ,h_w \mc T^\vee h_w^{-1}) \text{ corresponds to } (\mc L_w, \mc T_w) .
\end{equation}
We denote the rational series (in $\Rep (G^\epsilon)$) associated to $s_\beta \in 
G^{\vee \epsilon}$ by $\Rep_{s_\beta} (G^\epsilon)$. By the correspondence \eqref{eq:E.23}
we may write $\Rep_{s_\beta}(G^\epsilon) = \Rep_{W(\mc L,\beta)} (G^\epsilon)$. Lemma 
\ref{lem:E.6} shows how to write a geometric series as a direct sum of rational 
series, with $s_\beta$ as in \eqref{eq:E.23}:
\begin{equation}\label{eq:E.14}
\Rep_{W\mc L}(G^\epsilon) = \bigoplus_{\beta \in \mf B_{\mc L} / 
\mr{Ad}_\epsilon (\Omega_{\mc L})} \Rep_{W(\mc L,\beta)}(G^\epsilon) =
\bigoplus_{\beta \in \mf B_{\mc L} / 
\mr{Ad}_\epsilon (\Omega_{\mc L})} \Rep_{s_\beta}(G^\epsilon) ,
\end{equation}
Let $\mc G \to \mc G_c$ be a regular embedding of reductive $\F_q$-groups \cite[\S 1.7]{GeMa},
so $Z(\mc G_c)$ is connected and $\mc G_\der = \mc G_{1,\der}$. As is common, we will analyse
series for $G^\epsilon$ in terms of those for $G_c^\epsilon$. Let $\mc T_c \subset 
\mc G_c$ be the maximal torus with $\mc T_c \cap \mc G = \mc T$. We fix an extension $\mc L_c$
of $\mc L$ to a character sheaf on $\mc T_c^\epsilon$, and we let $s_c \in T_c^\epsilon$ be
the corresponding element. The canonical quotient map $\mc G_c^\vee \to \mc G^\vee$ sends
$\mc T_c^\vee \to \mc T^\vee$ and $s_c$ to $s$. The group $W_c = W(\mc G_c,\mc T_c)$ can
be identified with $W$. Since $\mc G_c$ has connected centre, $\mc H_c = Z_{\mc G_c}(s_c)$
is connected and
\[
W_{s_c} = W_{s_c}^\circ \cong W_{\mc L_c} = W_{\mc L_c}^\circ .
\]
The rational and geometric series for $G_c^\epsilon$ coincide, and by \cite{Lus2}:
\begin{equation}\label{eq:E.15}
\Rep_s (G^\epsilon) = \Res_{G^\epsilon}^{G_c^\epsilon} \big( \Rep_{s_c}(G_c^\epsilon) \big)
= \Res_{G^\epsilon}^{G_c^\epsilon} \big( \Rep_{W_{\mc L_c}}(G_c^\epsilon) \big) .
\end{equation}
Since $\mf B_{\mc L_c} = \{1\}$ and $\Omega_{\mc L_c} = \{1\}$, Theorem \ref{thm:E.1} for
$G_c^\epsilon$ simplifies to 
\begin{equation}\label{eq:E.16}
\Rep_{W \mc L_c}(G_c^\epsilon) \cong \Rep_1 (H_c^\epsilon) .
\end{equation}
The group $\mc H_c$ is generated by $\mc T_c$ and the root subgroups $U_\alpha$ with
$\alpha^\vee (s_c) = 1$. The latter condition depends only on $s$, so these are precisely
the roots in $\Phi_{\mc L}$ and $\mc H_c$ is the full preimage of $\mc H$ in $\mc G_c$.

\begin{thm}\label{thm:E.7}
Let $(\mc L,\mc T)$ be $\epsilon$-stable, corresponding to $(s,\mc T^\vee)$ with 
$s \in G^{\vee \epsilon}$. There is a canonical equivalence of categories 
\[
\Rep_s (G^\epsilon) = \Rep_{W(\mc L,1)} (G^\epsilon) \cong 
\Rep_1 (H^\epsilon )^{\Omega_{\mc L}^\epsilon} ,
\]
where $\Omega_{\mc L}^\epsilon \cong \pi_0 (Z_{\mc G^\vee}(s))^\epsilon$.
\end{thm}
\begin{proof}
We will use the functorial properties of the constructions of Lusztig and Yun. The functor
\[
\Res_{G^\epsilon}^{G_c^\epsilon} : \Rep_{W \mc L_c}(G_c^\epsilon) \to \Rep_{W \mc L}(G^\epsilon)
\]
corresponds to restriction of sheaves from $\mc G_c$ to $\mc G$ in \cite{LuYu}. Via 
\eqref{eq:E.16} and Theorem \ref{thm:E.1}, it becomes a functor
\begin{equation}\label{eq:E.18}
\Rep_1 (H_c^\epsilon) \longrightarrow
\bigoplus\nolimits_{\beta \in \mf B_{\mc L} / \mr{Ad}_\epsilon (\Omega_\mc L)} 
\Rep_1 \big( H^{\sigma_{\beta \epsilon}} \big)^{\Omega_{\mc L,\beta}} .
\end{equation}
From \cite[\S 12.10]{LuYu} one can see what \eqref{eq:E.18} does. Namely, first it restricts
representations of $H_c^\epsilon$ to $H^\epsilon$, and then it endows them with some
equivariant structure for $\Omega_{\mc L,\beta} = \mr{Stab}_{\Omega_{\mc L}}(\beta)$ with
$\beta = 1 \in \mf B_{\mc L}$. In particular the image of \eqref{eq:E.18} lies entirely in
the summand $\beta = 1$ of the codomain. Together with \eqref{eq:E.15}, it follows that the
image of $\Rep_s (G^\epsilon)$ under Theorem \ref{thm:E.1} is contained in 
$\Rep_1 (H^\epsilon)^{\Omega_{\mc L,1}}$.

Now we consider another $\epsilon$-fixed element in the geometric conjugacy class of $s$,
namely $s_\beta = h_w s h_w^{-1}$ from \eqref{eq:E.14} with 
$\beta \in \mf B_{\mc L} = W_{\mc L}^\circ \backslash W_{\mc L}$ represented by 
$w = w^\beta$ of minimal length in its class. We saw in \eqref{eq:E.20}
that it corresponds to an $\epsilon$-stable pair $(\mc L_w, \mc T_w) =
(g_w \mc L, g_w \mc T g_w^{-1})$. The above constructions can also be applied
to $(s_\beta,\mc L_w,\mc T_w)$, and they show that
\[
\Res_{G^\epsilon}^{G_c^\epsilon} : \Rep_{W \mc L_{w,1}}(G_c^\epsilon) \to 
\Rep_{(g_w W g_w^{-1}) \mc L_w}(G^\epsilon) = \Rep_{W \mc L}(G^\epsilon)
\]
corresponds to a functor 
\begin{equation}\label{eq:E.21}
\Rep_1 \big( (g_w H g_w^{-1})_c^\epsilon \big) \to 
\Rep_1 \big( (g_w H g_w^{-1})^\epsilon \big)^{\Omega_{\mc L_w ,1}} .
\end{equation}
When we restore the previous bookkeeping by conjugation with $g_w^{-1}$, the Frobenius
action $\epsilon$ is replaced by $\mr{Ad}(g_w^{-1} \epsilon (g_w)) \epsilon$, which 
equals $\mr{Ad}(\dot{w}^\beta) \circ \epsilon$ for some lift $\dot{w}^\beta \in 
N_{\mc G}(\mc T)$ of $w^\beta$. Following the conventions in \cite[\S 12.1]{LuYu} as 
mentioned after \eqref{eq:E.26}, \eqref{eq:E.21} becomes a functor
\[
\Rep_1 \big( H_c^{\sigma_{\beta \epsilon}} \big) \to 
\Rep_1 \big( H^{\sigma_{\beta \epsilon}} \big)^{\Omega_{\mc L,\beta}} .
\]
Thus Theorem \ref{thm:E.1} sends 
\begin{equation}\label{eq:E.24}
\Rep_{s_\beta}(G^\epsilon) = \Res_{G^\epsilon}^{G_c^\epsilon} \big( \Rep_{W \mc L_{w,1}}
(G_c^\epsilon) \big) \text{ to }
\Rep_1 \big( H^{\sigma_{\beta \epsilon}} \big)^{\Omega_{\mc L,\beta}}.
\end{equation}
If we take the direct sum over $\beta \in \mf B_{\mc L} / \mr{Ad}_\epsilon (\Omega_{\mc L})$
of the functors \eqref{eq:E.24}, then by \eqref{eq:E.14} we recover the equivalence from
Theorem \ref{thm:E.1}. Hence \eqref{eq:E.24} is an equivalence of categories for every
$\beta \in \mf B_{\mc L}$.

In the special case $\beta = 1$, we have 
\begin{equation}\label{eq:E.29}
\Omega_{\mc L,1} = \{ \omega \in \Omega_{\mc L} : \omega \epsilon (\omega)^{-1} \in 
W_{\mc L}^\circ \} = \{ \omega \in \Omega_{\mc L} : \omega \epsilon (\omega)^{-1} = 1 \}
= \Omega_{\mc L}^\epsilon .
\end{equation}
From \eqref{eq:E.17} we see that this group is canonically isomorphic to 
$\pi_0 (Z_{\mc G^\vee}(s))^\epsilon$.
\end{proof}

\section{Disconnected groups} 

We no longer require that $\mc G$ is connected. Instead, we assume the following throughout this
section:

\begin{cond}\label{cond:E.3}
$\mc G^\epsilon$ is a smooth $\F_q$-group scheme with reductive neutral component, such that
every connected component of $\mc G$ is $\epsilon$-stable.
\end{cond}

If $\mc G$ would have connected components that are not $\epsilon$-stable, then they would not 
contribute to $G^\epsilon$ and could be ignored for the study of $G^\epsilon$-representations. 
Hence the last part of Condition \ref{cond:E.3} is not a restriction, it merely serves to simplify 
the notations. Notice that $\pi_0 (\mc G)$ may be infinite.

Let $\pi_0^\epsilon (\mc G)$ be the group of connected components of $\mc G$. 
If $x \in \mc G$ represents an element of $\pi_0 (\mc G)$, then 
$x^{-1} \epsilon (x) \in \mc G^\circ$. By Lang's theorem there exist $g \in \mc G^\circ$ 
with $g^{-1} \epsilon (g) = x^{-1} \epsilon (x)$. Then $x g^{-1} \in x \mc G^\circ$ 
is fixed by $\epsilon$, so $x \mc G^\circ$ has $\F_q$-points.

For each $\omega \in \pi_0 (\mc G)$ we pick a representative $\dot \omega \in G^\epsilon = 
\mc G^\epsilon (\F_q)$. That gives a map
\[
\mr{Ad}(\dot \omega) : \Rep (G^{\circ \epsilon}) \to \Rep (G^{\circ \epsilon}).
\]
It depends on the choice of $\dot \omega$, but the map $\mr{Ad}(\ddot \omega)$ for another
representative $\ddot \omega$ differs only by $\mr{Ad}(\ddot{\omega} \dot{\omega}^{-1})$ with
$\ddot{\omega} \dot{\omega}^{-1} \in G^{\circ \epsilon}$. Hence $\mr{Ad}(\dot \omega)$ is
unique up to canonical isomorphisms. 

A $\pi_0 (\mc G)$-equivariant structure on $\pi^\circ \in \Rep (G^{\circ \epsilon})$
consists of a morphism $\pi (\dot \omega) : \pi^\circ \to \mr{Ad}(\dot \omega) \pi^\circ$ 
for each $\omega \in \pi_0 (\mc G)$. These must be multiplicative up to canonical 
isomorphisms, which means that
\[
\pi (\dot \omega_1) \pi (\dot \omega_2) = \pi^\circ (\dot \omega_1 \dot \omega_2
(\dot \omega_3)^{-1}) \pi (\dot \omega_3)
\]
whenever $\omega_1 \omega_2 = \omega_3 \in \pi_0 (\mc G)$.

\begin{lem}\label{lem:E.2}
The category $\Rep (G^{\circ \epsilon})^{\pi_0 (\mc G)}$ of 
$G^{\circ \epsilon}$-representations enriched with a $\pi_0 (\mc G)$-equivariant
structure is naturally equivalent with $\Rep (G^\epsilon)$.
\end{lem}
\begin{proof}
For any $G^\epsilon$-representation $\pi$, write $\pi^\circ = \Res^{G^\epsilon}_{G^{\circ \epsilon}}
\pi$. We obtain a $\pi_0 (\mc G)$-equivariant structure on $\pi^\circ$ by defining
the required morphism $\pi^\circ \to \mr{Ad}(\dot \omega) \pi^\circ$ as $\pi (\dot \omega)$, for any
representative $\dot \omega$ of $\omega \in \pi_0 (\mc G)$.

Conversely, suppose that $\pi^\circ \in \Rep (G^{\circ \epsilon})$ has a $\pi_0 (\mc G)
$-equivariant structure. Then we have a morphism of $G^{\circ \epsilon}$-representations
$\pi (\dot \omega) : \pi^\circ \to \mr{Ad}(\dot \omega) \pi^\circ$, for any representative 
$\dot \omega \in G^\epsilon$ of $\omega \in \pi_0 (\mc G)$. By the multiplicativity of
equivariant structures, these $\pi (\dot \omega)$ combine to a representation of 
\[
\bigcup\nolimits_{\omega \in \pi_0 (\mc G)} \, \omega G^{\circ \epsilon} = G^\epsilon 
\]
on the vector space underlying $\pi^\circ$. It follows from $\pi (1) = \mr{id}$ that 
$\pi (g) = \pi^\circ (g)$ for all $g \in G^{\circ \epsilon}$, so $\pi$ extends $\pi^\circ$.

It is clear that the functors $\pi \mapsto \pi^\circ$ and $\pi^\circ \mapsto \pi$ are mutually
inverse, so they define equivalences of categories.
\end{proof}

Let $\mc T$ be a maximal $\F_q$-torus in $\mc G^\circ$ and let $\mc T_0$ be a maximally split 
maximal $\F_q$-torus in $\mc G^\circ$. Let $\mc B$ and $\mc B_0$ be Borel $\F_q$-subgroup of 
$\mc G^\circ$ containing, respectively, $\mc T$ and $\mc T_0$. Although $\mc T$ may equal to
$\mc T_0$, we remain more flexible by not requiring that.

\begin{lem}\label{lem:E.9}
The group $G^\epsilon$ equals $G^{\circ \epsilon} N_{G^\epsilon}(T_0^\epsilon, B_0^\epsilon)$.
\end{lem}
\begin{proof}
For $x \in G^\epsilon$, $x \mc B_0 x^{-1}$ is an $\epsilon$-stable Borel subgroup of $\mc G^\circ$.
All such subgroups are $G^{0\epsilon}$-conjugate \cite[15.4.6.ii]{Spr}, so there exists 
$g \in G^{\circ \epsilon}$ such that $g x \mc B_0 x^{-1} g^{-1} = \mc B_0$. 

Let $\mc S$ be the maximal $\F_q$-split subtorus of $\mc T_0$. Since $\mc G^\circ$ is 
quasi-split over $\F_q$, the centralizer $Z_{\mc G^\circ}(\mc S)$ equals $\mc T_0$. 
As $gx \in N_{G^\epsilon}(B_0^\epsilon)$, $gx \mc S x^{-1} g^{-1}$ is a maximal $\F_q$-split 
torus in $\mc B_0$. All such tori are $B^\epsilon$-conjugate \cite[14.4.3]{Spr}, so we can 
find $b \in B^\epsilon$ such that  $bgx \mc S (bgx)^{-1} = \mc S$. Then
\[
bgx \mc T_0 (bgx)^{-1} = bgx Z_{\mc G^\circ}(\mc S) (bgx)^{-1} = Z_{\mc G^\circ} \big( bgx
\mc S (bgx)^{-1} \big) = Z_{\mc G^\circ} (\mc S) = \mc T_0 ,
\]
so $bgx \in N_{G^\epsilon}(B_0^\epsilon,T_0^\epsilon)$. It follows that $x \in G^{\circ \epsilon}
N_{G^\epsilon}(B_0^\epsilon,T_0^\epsilon)$.
\end{proof}

Lemma \ref{lem:E.9} implies that the natural map $N_{G^\epsilon}(T_0^\epsilon, B_0^\epsilon) / 
T_0^\epsilon \to \pi_0 (\mc G)$ is a group isomorphism. That yields split short exact sequences
\begin{equation}\label{eq:E.10}
\begin{array}{ccccccccc}
1 & \to & W^\circ = W(\mc G^\circ,\mc T) & \to & W = W (\mc G,\mc T) & \to & \pi_0 (\mc G) & \to & 1 ,\\
1 & \to & W(G^{\circ \epsilon},T_0^\epsilon) & \to & W (G^\epsilon,T_0^\epsilon) & 
\to & \pi_0 (\mc G) & \to & 1 . \\
\end{array}
\end{equation}
Condition \ref{cond:E.3} enables us to pick representatives $\dot \omega \in 
N_{G^\epsilon}(T_0^\epsilon, B_0^\epsilon)$ for $\omega \in \pi_0 (\mc G)$, unique up to $T_0^\epsilon$. 
The action of $\pi_0 (\mc G)$ on $\Rep (G^{\circ \epsilon})$ permutes the subcategories 
$\Rep_{W^\circ \mc L} (G^{\circ \epsilon})$ and stabilizes the subcategory
\begin{equation}\label{eq:E.8}
\Rep_{W \mc L}(G^{\circ \epsilon}) = \bigoplus_{w \in W^\circ \setminus W / W_{\mc L}}
\Rep_{W^\circ w \mc L} (G^{\circ \epsilon})
\end{equation}
generated by irreducible $G^{\circ\epsilon}$-representations with semisimple parameter in $W \mc L$. 
Indeed, for $\mc T = \mc T_0$ we see this with the representatives of $\pi_0 (\mc G)$ in
$N_{G^\epsilon}(T_0^\epsilon, B_0^\epsilon)$. For other $\mc T$ it follows because $(\mc T,\mc B)$
and $(\mc T_0,\mc B_0 )$ are $\mc G^\circ$-conjugate and $\Rep_{W^\circ \mc L} (G^{\circ \epsilon})$ 
depends only on $W^\circ \mc L$ up to $\mc G^\circ$-conjugacy. 

We define the geometric series $\Rep_{W \mc L}(G^\epsilon)$ as the category of 
$G^\epsilon$-representations whose restriction to $G^{\circ \epsilon}$ lies in 
$\Rep_{W \mc L}(G^{\circ \epsilon})$. Thus 
Lemma \ref{lem:E.2} provides a natural equivalence of categories
\begin{equation}\label{eq:E.7}
\Rep_{W \mc L}(G^\epsilon) \cong \Rep_{W \mc L}(G^{\circ \epsilon})^{\pi_0 (\mc G)} .
\end{equation}
The group $\pi_0 (\mc G) \cong W / W^\circ$ permutes the direct summands in \eqref{eq:E.8} 
transitively, which leads to a natural equivalence
\begin{equation}\label{eq:E.4}
\Rep_{W \mc L}(G^{\circ \epsilon})^{\pi_0 (\mc G)} \cong 
\Rep_{W^\circ \mc L}(G^{\circ \epsilon})^{\pi_0 (\mc G)_{W^\circ \mc L}} . 
\end{equation}
We define $\Omega_{\mc L} = \mr{Stab}_{W_{\mc L}}(\Phi_{\mc L}^+) \cong W_{\mc L} / W_{\mc L}^\circ$ 
as before, and we let $\Omega_{\mc L}^\circ \cong (W^\circ)_{\mc L} / W_{\mc L}^\circ$ 
be its version for $\mc G^\circ$. Then
\begin{equation}\label{eq:E.5}
\pi_0 (\mc G)_{W^\circ \mc L} \cong \Omega_{\mc L} / \Omega_{\mc L}^\circ .
\end{equation}
Let $\mf B_{\mc L}^\circ$ be $\mf B_{\mc L}$ for $\mc G^\circ$. With Theorem \ref{thm:E.1} we can
transfer the action of $\pi_0 (\mc G)_{W^\circ \mc L}$ on $\Rep_{W^\circ \mc L} (G^{\circ \epsilon})$
to an action on 
\begin{equation}\label{eq:E.9}
\Big( \bigoplus\nolimits_{\beta \in \mf B_{\mc L}^\circ} 
\Rep_1 \big( H^{\sigma_{\beta \epsilon}} \big) \Big)^{\Omega_{\mc L}^\circ} ,
\end{equation}
unique up to canonical isomorphisms. Let us make that more explicit. The action of 
$\pi_0 (\mc G)_{W^\circ \mc L}$ on $\bigoplus\nolimits_{\beta \in \mf B_{\mc L}^\circ} \Rep_1 
(H^{\sigma_{\beta \epsilon}})$ can be described in the same way as the action of $\Omega_{\mc L}^\circ$ 
in \eqref{eq:E.28}, now only with representatives $n(\omega,\beta)$ from $N_{\mc G}(\mc T,\mc B)$. 
The effect on a morphism $\lambda_{\dot w} : \pi_\beta \to \mr{Ad}(\dot{w}) \pi_\beta$, from an 
$\Omega_{\mc L}^\circ$-equivariant structure $\lambda$ on $\pi = \bigoplus_\beta \pi_\beta$, 
is the morphism
\[
\mr{Ad}(n(\omega,\beta)) \pi_\beta \to \mr{Ad}(\dot w) \mr{Ad}(n(\omega,\beta)) \pi_\beta
\]
given by 
\[
\lambda_{n(\omega,\beta)^{-1} \dot{w} n(\omega,\beta)} : \pi_\beta \to 
\mr{Ad}(n(\omega,\beta)^{-1}) \mr{Ad}(\dot{w}) \mr{Ad}( n(w,\beta)) \pi_\beta .
\]
Via this transfer of actions, every $\pi_0 (\mc G)_{W^\circ \mc L}$-equivariant
structure on \eqref{eq:E.9} can be identified with one $\pi_0 (\mc G)_{W^\circ \mc L}$-equivariant 
structure on $\Rep_{W^\circ \mc L} (G^{\circ \epsilon})$.

The group $\Omega_{\mc L}$ acts on $\bigoplus\nolimits_{\beta \in \mf B_{\mc L}^\circ} 
\Rep_1 (H^{\sigma_{\beta \epsilon}})$, by the same constructions as in the case of a connected 
$\mc G$. Again this action is well-defined up to canonical isomorphisms, and there is a notion of
$\Omega_{\mc L}$-equivariant structures on $\bigoplus\nolimits_{\beta \in \mf B_{\mc L}^\circ} 
\Rep_1 (H^{\sigma_{\beta \epsilon}})$.

\begin{thm}\label{thm:E.4}
Suppose that $\mc G^\epsilon$ satisfies Condition \ref{cond:E.3}, and let $\mc L$ be a rank one 
character sheaf on $\mc T$, such that $W^\circ \mc L$ 
is $\epsilon$-stable. There exists a canonical equivalence of categories
\[
\Rep_{W \mc L} (G^\epsilon) \cong \Big( \bigoplus\nolimits_{\beta \in \mf B_{\mc L}^\circ} 
\Rep_1 \big( H^{\sigma_{\beta \epsilon}} \big) \Big)^{\Omega_{\mc L}}. 
\]
\end{thm}
\begin{proof}
By Lemma \ref{lem:E.2}, \eqref{eq:E.7}, \eqref{eq:E.4} and Theorem \ref{thm:E.1}, 
there are canonical equivalences of categories
\begin{equation}\label{eq:E.6}
\Rep_{W \mc L} (G^\epsilon) \cong \Rep_{W^\circ \mc L} (G^{\circ \epsilon})^{\pi_0 (\mc G)_{W^\circ 
\mc L}} \cong \Big( \big( \bigoplus\nolimits_{\beta \in \mf B_{\mc L}^\circ} 
\Rep_1 (H^{\sigma_{\beta \epsilon}}) \big)^{\Omega_{\mc L}^\circ} \Big)^{\pi_0 (\mc G)_{W^\circ \mc L}} \!.
\end{equation}
In view of \eqref{eq:E.5}, a $\pi_0 (\mc G)_{W^\circ \mc L}$-equivariant structure on top of a 
$\Omega_{\mc L}^\circ$-equivariant structure contains precisely the same information as an
$\Omega_{\mc L}$-equivariant structure. Hence the right hand side of \eqref{eq:E.6} is naturally
equivalent to $\bigoplus\nolimits_{\beta \in \mf B_{\mc L}^\circ} 
\Rep_1 (H^{\sigma_{\beta \epsilon}})$ enriched with $\Omega_{\mc L}$-equivariant structures.
\end{proof}

We turn to rational series for $G^\epsilon$. First we have to define them properly.
By \eqref{eq:E.23}, we can parametrize the rational series in $\Rep (G^{\circ \epsilon})$
by $W^\circ$-orbits of pairs $(\mc L,\beta)$. If we set that up with $\mc T_0$, \eqref{eq:E.10}
shows that the groups $\pi_0 (\mc G) \cong N_{G^\epsilon}(T_0^\epsilon,B_0^\epsilon)/T_0^\epsilon$
and $W(\mc G,\mc T_0) \cong W$ and act naturally on the set of such parameters. Then the 
definition of $\Rep_{W^\circ (\mc L,\beta)}(G^{\circ \epsilon})$ in terms of generators obtained
by Deligne--Lusztig induction shows that $\pi_0 (\mc G)$ and $W$ and permute the
various rational series according to the actions on their parameters. 
It follows that the category
\[
\Rep_{W(\mc L,\beta)} (G^{\circ \epsilon}) := \bigoplus\nolimits_{\gamma \in W^\circ W_{\mc L,\beta}
\backslash W} \Rep_{W^\circ \gamma \mc L, \gamma \beta \epsilon (\gamma)^{-1}} (G^{\circ \epsilon})
\]
is stable under the action of $G^\epsilon$ on $\Rep (G^\circ)$. Via conjugation in $\mc G^\circ$,
these conside\-ra\-tions also apply to any $\epsilon$-stable maximal torus $\mc T$ instead of $\mc T_0$.

We define $\Rep_{W(\mc L,\beta)} (G^\epsilon)$ as the ca\-te\-gory of 
$G^\epsilon$-representations whose restriction to $G^{\circ \epsilon}$ lies in 
$\Rep_{W(\mc L,\beta)} (G^{\circ \epsilon})$. When $W^\circ (\mc L,\beta)$ corresponds to 
$\mr{Ad}(G^{\circ \vee \epsilon}) s_\beta$, we write 
\[
\Rep_{W(\mc L,\beta)} (G^\epsilon) = \Rep_{s_\beta}(G^\epsilon). 
\]
We call this the rational series of $G^\epsilon$ associated to $s_\beta$ or to $W(\mc L,\beta)$. 
It actually corresponds to a set of semisimple elements in $G^{\circ \vee \epsilon}$ larger than 
the rational conjugacy class of $s_\beta$, but it is difficult to express that precisely because 
we do not have a group dual to $\mc G$.

As the analysis of $\Rep_{W(\mc L,\beta)} (G^\epsilon)$ proceeds like for $G^{\circ \epsilon}$, 
we will carry it out more briefly. Like in \eqref{eq:E.7} and \eqref{eq:E.4} we have
\begin{equation}\label{eq:E.25}
\Rep_{W (\mc L,\beta)}(G^\epsilon) \cong \Rep_{W (\mc L,\beta)}(G^{\circ \epsilon})^{\pi_0 (\mc G)} 
\cong \Rep_{W^\circ (\mc L,\beta)}(G^{\circ \epsilon})^{\pi_0 (\mc G)_{W^\circ (\mc L,\beta)}} . 
\end{equation}
There are isomorphisms
\begin{equation}\label{eq:E.27}
\pi_0 (\mc G)_{W^\circ (\mc L,\beta)} \cong W_{\mc L,\beta} / (W^\circ)_{\mc L,\beta} =
(W_{\mc L}^\circ \rtimes \Omega_{\mc L,\beta}) / (W_{\mc L}^\circ \rtimes \Omega^\circ_{\mc L,\beta})
\cong \Omega_{\mc L, \beta} / \Omega_{\mc L,\beta}^\circ .
\end{equation}
By Lemma \ref{lem:E.5} it suffices to consider the cases where $\mc L$ is $\epsilon$-stable and 
$\beta = 1$. 

\begin{thm}\label{thm:E.8}
Suppose that $\mc G^\epsilon$ satisfies Condition \ref{cond:E.3}. Let $\mc L$ be an 
$\epsilon$-stable rank one character sheaf on $\mc T$ and suppose that $W^\circ (\mc L,1)$ 
corresponds to $s \in G^{\circ \vee \epsilon}$ via \eqref{eq:E.23}. There exists a 
canonical equivalence of categories 
\[
\Rep_s (G^\epsilon) = \Rep_{W(\mc L,1)} (G^\epsilon) \cong 
\Rep_1 (H^\epsilon )^{\Omega_{\mc L}^\epsilon} .
\] 
\end{thm}
\begin{proof}
This is entirely analogous to Theorem \ref{thm:E.4}. By Lemma \ref{lem:E.2}, \eqref{eq:E.25} 
and Theorem \ref{thm:E.7}, there are canonical equivalences of categories
\[
\Rep_{W (\mc L,1)} (G^\epsilon) \cong \Rep_{W^\circ (\mc L,1)} (G^{\circ \epsilon})^{\pi_0 
(\mc G)_{W^\circ (\mc L,1)}} \cong \Big( \Rep_1 (H^\epsilon)^{\Omega_{\mc L,1}^\circ} 
\Big)^{\pi_0 (\mc G)_{W^\circ (\mc L,1)}} .
\]
In the rightmost term, the two equivariant structures combine to one 
$\Omega_{\mc L,1}$-equivariant structure, by \eqref{eq:E.27}.
We recall from \eqref{eq:E.29} that $\Omega_{\mc L,1} = \Omega_{\mc L}^\epsilon$.
\end{proof}

\textbf{Acknowledgement.}
The author thanks Tasho Kaletha for many enlightening discussions on this and
related topics.

\end{document}